OpenStudy (anonymous):

125 ^ 9x-2 = 150 can anyone help me figure out how to solve this step by step ?

hartnn (hartnn):

is that \(1259x-2=150\) ? start by adding 2 on both sides...

hartnn (hartnn):

but i don't think thats your question, you could've easily solved that

OpenStudy (anonymous):

sorry , i mis wrote it , 9x - 2 is an exponent of 125

OpenStudy (anonymous):

i think i have to use log but im not any good at logarithm /:

hartnn (hartnn):

\(125 ^{9x-2} = 150\) taking log on both sides

hartnn (hartnn):

use the property that \(\large \log a^b = b \log a \) on left side

OpenStudy (anonymous):

9x-2 log ( 125 ) = log ( 150 ) is that right so far ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes it is but don't forget the brackets (9x-2) log 125 = log 150

hartnn (hartnn):

i think now you can isolate the 'x'

OpenStudy (anonymous):

but couldnt it also be log base 125 ^ 150 = 9x-2 ?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, it could be you have used change of base formula

hartnn (hartnn):

\(9x-2 = \log_{125}150\) add 2 and then divide by 9

OpenStudy (anonymous):

so log 150 / log 125 = 9x - 2 ?

OpenStudy (anonymous):

i got .337 /.\ is that right ?

hartnn (hartnn):

yeah i get that too weird answer :O

OpenStudy (anonymous):

thank you soo muuch ! (:

hartnn (hartnn):

welcome ^_^